# Can ChatGPT Detect DeepFakes? A Study of Using Multimodal Large Language Models for Media Forensics

### v3

@@ 3 Methodology
  • Prompt #3: Tell me the probability of this image being real. Answer a probability score between 0 and 100.
  • Prompt #4: Tell me the probability of this image being AI-generated. Answer a probability score between 0 and 100.
@@
  • Prompt #6: Tell me if synthesis artifacts are in the face. Must return with 1) yes or no only; 2) if yes, explain where the artifacts exist by answering in [region, artifacts] form.
@@ 4.1 Qualitative and Quantitative Results
On the other hand, we note that most errors of GPT4V occur on real images – per Table 3, the classification accuracies on real images are around $50\%$, drastically different from those of AI-generated images, which is above $90\%$. Some intuitions can be obtained when we examine the real face images for these error cases, as shown in Fig. 4. These cases include semantic features unusual for “typical” face images. For instance, different age group (baby in the first case) or unique hair color (second case) to style (third case). This suggests that the semantic abnormality identified by GPT4V may not be specific to DeepFake faces. This problem may be solved by refining the model. In contrast, the Gemini model achieves a classification accuracy of $83.3\%$ on real images, dropping to around $50\%$ on generated faces. The examples in Fig. 5 show that the Gemini model’s response lacks rationality in analyzing the synthesis artifacts.
@@ 5 Conclusion
In this study, we investigate the potential of leveraging multimodal LLMs for tasks related to media forensics. Our research will broaden the application of multimodal LLMs to include a wider array of media forms, particularly focusing on video analysis. Rather than simply applying image-based detection techniques to video frames, a more integrated approach would involve direct video content processing. Additionally, we aim to enhance the detection of text-image mis-contextualization [15], where images and text are misleadingly paired to spread false information. Future endeavors will also explore developing more sophisticated prompting strategies and integrating these models with conventional signal or data-driven detection techniques.

### v4

@@ 3 Methodology
  • Prompt #3: Tell me the probability of this image being AI-generated. Answer a probability score between 0 and 100.
  • Prompt #4: Tell me the probability of this image being real. Answer a probability score between 0 and 100.
@@
  • Prompt #6: Tell me if there are synthesis artifacts in the face or not. Must return with 1) yes or no only; 2) if yes, explain where the artifacts exist by answering in [region, artifacts] form.
@@ 4.1 Qualitative and Quantitative Results
On the other hand, we note that most errors of GPT4V occur on detecting real images – per Table 3, the classification accuracies on real images are around $50\%$, drastically different from those of AI-generated images, which is above $90\%$. Some intuitions can be obtained when we examine the real face images for these error cases, as shown in Fig. 4. These cases include semantic features unusual for “typical” face images. For instance, different age group (baby in the first case) or unique hair color (second case) to style (third case). This suggests that the semantic abnormality identified by GPT4V may not be specific to DeepFake faces. This problem may be solved by refining the model. In contrast, the Gemini model achieves a classification accuracy of $83.3\%$ on real images, dropping to around $50\%$ on generated faces. The examples in Fig. 5 show that the Gemini model’s response lacks rationality in analyzing the synthesis artifacts.
@@ 5 Conclusion
In this study, we investigate the potential of leveraging multimodal LLMs for tasks related to media forensics. Our future research will broaden the application of multimodal LLMs to include a wider array of media forms, particularly focusing on video analysis. Rather than simply applying image-based detection techniques to video frames, a more integrated approach would involve direct video content processing. Additionally, we aim to enhance the detection of text-image mis-contextualization [15], where images and text are misleadingly paired to spread false information. Future endeavors will also explore developing more sophisticated prompting strategies and integrating these models with conventional signal or data-driven detection techniques.